AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Sat Apr 5 00:42:50 2025
Author
*
:
Subject
*
:
09.00 CARME IE514 5.1e-11mbar 09.15 IE514 not working - one of cables disconnected? DSSSD #1 bias cable changed from SHV - Lemo 00.250 to SHV - 2x Lemo 00.250 DSSSD #1 & #2 LK1 removed from n+n Ohmic side bias FEE64s aida11 HDMI cabling re-seated at MACB 10.26 DSSSD bias OK - attachment 1 Ambient temperature +16.2 deg C 10.30 ADC data item stats OK - attachment 2 aida06 & aida15 > 100k, all others c. 3k consistent with 50Hz pulser 10.34 per FEE64 Rate spectra - attachment 3 OK except for known issues aida05, aida07 and aida16 - all others OK per p+n FEE64s 1.8.L spectra - attachment 4 pulser peak width 52 ch FWHM aida11- double peaking DSSSD #1 & #2 per n+n FEE64s 1.8.L spectra - attachment 7 pulser peak width 77 ch FWHM aida04 - more uniform noise across all DSSSDs 10.42 per FEE64 1.8.W spectra - 20us FSR - attachments 5-6 11:00 LK2 removed from aida01-04. Configuration for all adaptor cards/FEE modules is now the same. No other change to the setup. 12:00 FEEs power on, bias to 150 V. No change in bias observed. 12.02 ADC data item stats OK aida06 & aida15 > 100k, all others c. 3k consistent with 50Hz pulser 12.07 per FEE64 Rate spectra OK except for known issues aida05, aida07 and aida16 - all others OK 12:09 per p+n FEE64s 1.8.L spectra We now observe a large peak around chn 30,000 on all FEEs. Perhaps an effect of some change in the ring? pulser peak width 53 ch FWHM aida11- still observe double peaking DSSSD #1 & #2 per n+n FEE64s 1.8.L spectra Lots of events at ~chn 65,000 - double peaking now observed on aida03 and aida04 pulser peak width 93 ch FWHM aida14 12.20 per FEE64 1.8.W spectra - 20us FSR Perhaps bias was off while acquiring and spectra was not cleared properly. Re-do tests 13:40 ADC data item stats OK - attachment 7 aida06 & aida15 > 100k, all others c. 3k consistent with 50Hz pulser per FEE64 Rate spectra - attachment 8 OK except for known issues aida05, aida07 and aida16 - all others OK per p+n FEE64s 1.8.L spectra - attachment 9 pulser peak width 55 ch FWHM aida11- double peaking DSSSD #1 & #2 per n+n FEE64s 1.8.L spectra - attachment 10 pulser peak width 65 ch FWHM aida14 - more uniform noise across all DSSSDs per FEE64 1.8.W spectra - 20us FSR - attachments 11-12 With help from NH updated firmware on all MACB modules to version 4/20 - see https://elog.ph.ed.ac.uk/CARME/482 Timestamps for all FEE modules now correct - attachment 13 MAC address of aida07 was found to be 41-b4-16 not 41-d4-16 as was labelled on the FEE module re-labelled FEE64 dhcpd.conf and MIDAS startup.tcl updated, dhcpd and MIDAS 8015 server restarted - aida07 is now working OK CAEN N1419ET internal jumper fitted - outputs are no longer floating 15:20 ADC data item stats OK aida06 & aida15 > 100k, all others c. 3k consistent with 50Hz pulser - attachment 14 per FEE64 Rate spectra OK except for known issues aida05, and aida16 - all others OK - attachment 15 per p+n FEE64s 1.8.L spectra - pulser peak width 50 ch FWHM aida11- double peaking DSSSD #1 & #2 - attachment 16 per n+n FEE64s 1.8.L spectra pulser peak width 67 ch FWHM aida14 - more uniform noise across all DSSSDs per FEE64 1.8.W spectra - 20us FSR - aida14 missing from waveforms - attachments 17,18 Swapped the test inputs so that +ve is now from the inverter and -ve direct from pulser. - attachments 19,20. Observed reduced noise for when plugged directly into the pulser than when connected from the inverter. Lowered threshold from 0x64 (1 MeV) to 0x20 (320 keV) - attachment 21, 22 - Observed increase in rates, in particular for the DSSD #1,2 pn side. Conclusion -> increased pulser width for DSSD #1,2 is real and not from the pulser aida01 lemo to water pipe (chamber ground) connection installed ADC data item stats OK aida06 & aida15 > 100k, all others c. 3k consistent with 50Hz pulser per FEE64 Rate spectra OK except for known issues aida16 - all others OK per p+n FEE64s 1.8.L spectra - pulser peak width 48 ch FWHM aida11- double peaking DSSSD #1 & #2 - attachment 23 per n+n FEE64s 1.8.L spectra pulser peak width 67 ch FWHM aida14 - attachment 24 per FEE64 1.8.W spectra - 20us FSR - slightly reduced 100 kHz on pn waveforms - attachments 25,26 Disconnected aida06 and replaced with the spare FEE module (mac address: 42:0d:16). Power cycled and updated config files. ADC data item stats OK aida15 > 100k, all others including aida06 c. 3k consistent with 50Hz pulser - attachment 27 per FEE64 Rate spectra OK except for known issues aida16 - all others OK - attachment 28 per p+n FEE64s 1.8.L spectra - pulser peak width 48 ch FWHM aida11- no longer observe double peaking DSSSD #1 & #2 but the pulser peaks are still broader compared to DSSD #3,4 - attachment 29 per n+n FEE64s 1.8.L spectra pulser peak width 55 ch FWHM aida14 - attachment 30 per FEE64 1.8.W spectra - 20us FSR - greatly reduced 100 kHz on pn waveforms, dodgy FEE previously aida06 must@ve had some impact on DSSD#1,2 - attachments 31,32
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown